Goodbye, Inc., recently issued new securities to finance a new TV show. The project cost $14.7 million, and the company paid $795,000 in flotation costs. In addition, the equity issued had a flotation cost of 7.7 percent of the amount raised, whereas the debt issued had a flotation cost of 3.7 percent of the amount raised.

If the company issued new securities in the same proportion as its target capital structure, what is the company’s target debt–equity ratio? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1

Let x be Equity and y be Debt in millions

Hence,
x+y=14.7
x*7.7%+y*3.7%=0.795

=>x=6.2775 million and
y=8.4225 million

Target Debt Equity Ratio=8.4225/6.2775=1.341696535